# Sequential Quadratic Sum-of-squares Programming for Nonlinear Control Systems

### v1

@@ 9  Update Hessian approximation
where $H_{k}:\mathbb{R}_{2d}[x]^{n}\to(\mathbb{R}_{2d}[x]^{n})^{*}$ is an approximation of the Hessian of the Lagrangian in (2). Some conic solvers [38, 22] can handle a quadratic cost directly, otherwise one rewrites the quadratic cost into a second-order cone constraint using the epigraph reformulation [7].
@@
where $\xi_{+}$, $\lambda_{+}$ and $\mu_{+}$ denote the optimal solution, corresponding Lagrange multipliers and normal vector of (3), respectively. This approach is also known as quasi-Newton method [25] and it exhibits fast local convergence with a well-chosen initial guess; however, it often requires a globalization technique for robust performance from arbitrary starting points. We are going to utilize a line search to compute the new candidate solution as
$\xi^{k+1}$ $=\xi^{k}$ $+\alpha^{k}\underbrace{(\xi_{+}-\xi^{k})}_{\omega^{k}}$ (5a)
@@
where $\alpha^{k}\in(0,1]$ is the chosen step length at iteration $k$. Approximation methods for the Hessian are discussed in Subsection IV-C1.
III-B Local Convergence Analysis
Should we explain, why we only consider the best case, i.e., neglecting any kind of globalization? How is without globalization the ‘best’ case? We should state though, preferably already in the introduction, that the proof by Wächter and Biegler considers active sets methods and cannot be applied to SOS; and that we leave the global convergence proof to future work.
@@ III-B Local Convergence Analysis
The convergence properties of set-valued Newton methods are a long studied subject of variational analysis citations. Local convergence can usually be established provided that the inverse of $(\Phi(\xi^{k},H_{k})+\mathcal{N})$ in (6) is well defined and the approximation $\Phi(\xi^{k},H_{k})$ is ‘sufficiently good’ – the better the approximation, the faster the rate of convergence. In particular, we make the following assumptions:
Assumption 1.
The exist constants $\zeta,\ell_{k}>0$ such that
@@
    the inverse $(\Phi(\xi^{k},H_{k})+\mathcal{N})^{-1}$ is locally single-valued around $(\xi_{\star},\lambda_{\star})$ and locally Lipschitz continuous at 0 with constant $\zeta$ uniformly for $\xi^{k}$ around $\xi_{\star}$ and all $H_{k}$;
@@
    the mapping $\varphi(\xi^{k},\lambda^{k})+\Phi(\xi^{k},H_{k})(\xi^{\prime}-\xi^{k},\lambda^{\prime}-\lambda^{k})$ is locally Lipschitz continuous relative to $(\xi^{k},\lambda^{k})$ at $(\xi_{\star},\lambda_{\star})$ with constant $\ell_{k}$ uniformly for $(\xi^{\prime},\lambda^{\prime})$ around $(\xi_{\star},\lambda_{\star})$ and all $H_{k}$.
Part (a) of Assumption 1 is called strong regularity of the set-valued mapping $(\Phi(\xi^{k},H_{k})+\mathcal{N})$ and corresponds to unique primal-dual solutions to the convex problem (3) that are continuous under first-order perturbations e.g., [15]. Regularity can be enforced by choosing a positive definite approximation $H_{k}$. The Lipschitz constant in part (b) measures how good the linear operator $\Phi(\xi^{k},H_{k})$ describes the change of $\varphi(\xi^{k},\lambda^{k})$ as $(\xi^{k},\lambda^{k})\to(\xi_{\star},\lambda_{\star})$ and can be thought of as generalization of an error in the derivative of $\varphi$. In particular, Assumption 1(b) is satisfied with $\ell_{k}=0$ if $H_{k}$ is the second-order derivative $\nabla_{\xi\xi}^{2}[f(\xi_{\star})-\langle\lambda_{\star},g(\xi_{\star})\rangle]$.
Theorem 2.
@@
Proof.
By virtue of [16, Theorem 8.5] with $h:(\xi,\lambda)\mapsto\varphi(\xi_{\star},\lambda_{\star})+\Phi(\xi_{\star},H_{k})(\xi-\xi_{\star},\lambda-\lambda_{\star})$, the sequence $\{(\xi^{k},\lambda^{k})\}_{k\in\mathbb{N}}$ generated by (6) is unique and satisfies
$\|(\xi^{k+1},\lambda^{k+1})-(\xi_{\star},\lambda_{\star})\|\leq\zeta\ell_{k}\|(\xi^{k},\lambda^{k})-(\xi_{\star},\lambda_{\star})\|$
for all $k\in\mathbb{N}$, provided that $(\xi^{k},\lambda^{k})$ is close to $(\xi_{\star},\lambda_{\star})$; this implies (super) linear convergence if $\zeta\ell_{k}<1$ (resp., $\zeta\ell_{k}\searrow 0$), and quadratic convergence if $\zeta\ell_{k}\equiv 0$. ∎
III-C Line-search with Filter
This subsection provides the core components of the filter line-search algorithm, which is primarily based on [6, p. 120]. We outline the core parts to make this paper self-contained. The filter line-search algorithm is summarized in Algorithm 2. In general three major steps are conducted: 1) check filter acceptance, 2) check sufficient decrease in cost and 3) augment filter, if necessary.
III-C1 Filter Acceptance
Unlike merit functions, that combine cost $f$ and constraint violation $\theta$ in a single objective, a filter tries to deal with these two conflicting goals in a bi-objective optimization manner. No penalty parameter must be estimated, as is the case for penalty functions or augmented Lagrangian. The filter consists of $l$ so called dominant pairs $\{f(\xi^{l}),\theta(\xi^{l})\in\mathcal{F}$ which form a forbidden region. In a first step, the current trial point is checked for filter acceptance, that is, checking if the trial point lies not in the forbidden region and makes sufficient progress regarding constraint violation. A new trial point, usually the result of (5) for the candidate step length $\alpha$, is accepted to the filter $\mathcal{F}$ if
$f(\xi^{k+1})<f(\xi^{l})\quad\text{or}\\ \quad\theta(\xi^{k+1})<\theta(\xi^{l})\quad\forall(f(\xi^{l}),\theta(\xi^{l}))\in\mathcal{F}$ (7)
Condition (7) alone only provides sufficient reduction of the constraint violation but not for improvement of the cost function. Therefore, additional sufficient decrease conditions for the cost function are checked for an acceptable point, as outlined next. If the candidate fails or violates the first condition (7), a second-order correction may be applied, or the step length is reduced.
III-C2 Sufficient Decrease Conditions
@@
where $s_{\phi},s_{\theta}\in(0,1)$ and $\delta\in(0,1)$ are small constants. This condition ensures a sufficiently large progress for the cost function compared to the current constraint violation [6]. Equations (8) and (9) are also referred as f-type switching condition.
In case the f-type switching is fulfilled and additionally if $\theta(\xi^{k+1})<\theta_{\text{min}}$ [51] then the Amijo condition is checked
@@
If no reduction in cost is possible (according to f-type) it is checked if either the constraint violation or cost is smaller than the previous iterate with a small envelope
$\theta(\xi^{k+1})\leq\theta(\xi^{k})(1-\gamma_{\theta})\text{ or }\phi(\xi^{k+1})\leq\phi(\xi^{k})-\gamma_{f}\theta(\xi^{k})$ (11)
@@
III-C3 Filter Augmentation
Unlike [20], the filter is only augmented if, and only if, any of the sufficient decrease conditions (8), (9) or (10) do not hold. This stricter augmentation rule avoids cycling, that is, the iterates alternately improve one of the two measures, but deteriorate the other [50].
III-C4 Full Algorithm
Algorithm 2 outlines the filter line-search procedure (cf.  [6]). To avoid the Maratos effect [36], we additionally add a second-order correction (details in Section IV-C2). If the second-order correction is able to find an acceptable trial point, the line-search is terminated, otherwise, the step length is reduced. In case the step-length falls below the minimum step-length $\alpha_{\text{min}}$ a feasibility restoration phase is invoked. The primary goal of the feasibility restoration phase is to minimize the constraint violation, such that the trial point is acceptable to the filter.
@@ 32   Augment filter
Algorithm 2 Filter Line-search with second-order correction (SOC)
III-D Required SOS Modification
The filter frequently needs to check progress in cost and feasibility. While checking progress in cost is a simple function evaluation, evaluating the constraint violation (feasibility) for SOS constraints is more complicated. In parameter optimization, constraint violation can be assessed by evaluating the constraints at the current iterate and use some norm, e.g., $\theta(\xi^{k}):=\|g(\xi^{k})\|_{\infty}$. For conic optimization, the constraint violation check involves a projection onto the corresponding cone (cf. [7]). Although some cones possess relatively simple (analytical) projection operators, this is not the case for the SOS cone. Since the constraint violation is frequently evaluated, we are interested in precise yet efficient methods for this task. Similarly, a dedicated feasibility restoration,First mention? which primary goal is to reduce constraint violation, must be designed and implemented for SOS.
IV Practical Implementation
[…]
@@ 35     Compute Hessian approximation
IV-A Constraint Violation Estimation
@@
We assume $n_{g}$ nonlinear constraints. The constraint violation $\theta(\xi^{k})$ for the solution $\xi^{k}$ indicates whether the constraints $g(\xi^{k})\in\Sigma[x]^{n_{g}}$ are satisfied ($\theta(\xi^{k})=0$) or, if not, measures how far ($\theta(\xi^{k})>0$) the solution is from satisfying the constraints. Take $p_{k}:=g(\xi^{k})$. Besides, the last statement is simply not true.
IV-A1 SOS Projection
@@
IV-A2 Signed Distance
In [14, Sec. 4.3], a signed-distance metric for generalized cones is introduced. It utilizes a scalarization function [27] to map $p_{k}$ to a finite signed distance between $p$ and the boundary of $\Sigma[x]^{n_{g}}$. The signed-distance metric is defined as
@@ IV-A2 Signed Distance
$\theta(x):=\begin{cases}0&\text{if }\mathrm{max}(r^{*})\leq\epsilon\\ \mathrm{max}(r^{*})&\text{otherwise}.\end{cases}$ (15)
@@
IV-A3 Pseudo-Projection via Sampling
Verifying that $p_{k}(x)\in\Sigma[x]^{n_{g}}$ can be replaced by checking the necessary condition $p_{k}(x)\geq 0$ on a finite set of sample points $x^{(i)}\in\mathbb{R}^{n}$, $i=1,\dots,M$, and nonnegativity is checked element-wise. Samples are drawn from a bounded hypercube.
@@ Discussion
Figure 1 compares the computation times of these methods for increasing numbers $n$ of indeterminate variables, using 10 randomly generated SOS polynomials of degree six. The projection method scales poorly, becoming impractical for larger $n$. In contrast, the signed distance remains efficient, while sampling with 1,000 points is fastest but least accurate.Should we report false-positives for Sampling? Based on this comparison, we implement the signed distance method for its strong balance of accuracy and efficiency.
@@
IV-B Feasibility Restoration
Small step lengths can cause the solver to stall or fail. If $\alpha^{k}<\alpha_{\text{min}}$ or infeasibility is detected in (3), Alg. 3 enters the feasibility restoration phase (lines 10 and 21 in Alg. 3) to maintain robustness [36, p.439]. This phase aims to reduce the constraint violation and provide a solution that is acceptable to the filter of the original problem.
Remark 1.
There is no guarantee that the restoration finds a solution acceptable to the filter. If it does not, we want it to converge to a local (infeasible) point. This information might be used by the user to take corrective action on the problem formulation (cf. [51]).
The feasibility restoration might be invoked after certain progress has already been made towards a local solution, but the constraint violation is still too high. In that case, we want the restoration not to deviate too much from the current, potentially close to (local) optimal solution. To achieve this, we introduce a weighted regularization term based on the current constraint violation [51, 49].

### v2

@@ 9  Update Hessian approximation
where $H_{k}:\mathbb{R}_{2d}[x]^{n}\to(\mathbb{R}_{2d}[x]^{n})^{*}$ is an approximation of the Hessian of the Lagrangian in (2). Some conic solvers scs ; Clarabel_2024 can handle a quadratic cost directly, otherwise one rewrites the quadratic cost into a second-order cone constraint using the epigraph reformulation Boyd2005 . In our implementation, we rely on existing SDP solvers.
@@
where $\xi_{+}$, $\lambda_{+}$ and $\mu_{+}$ denote the optimal solution, corresponding Lagrange multipliers, and normal vector of (4), respectively. The (primal) search direction in Alg. 1 is then given as $\omega^{k}=\xi_{+}-\xi^{k}$. This approach is also known as quasi-Newton method Izmailov2014 and it exhibits fast local convergence with a well-chosen initial guess; however, it often requires a globalization technique for robust performance from arbitrary starting points. We are going to utilize a line search to compute the new candidate solution as
$\xi^{k+1}$ $=\xi^{k}$ $+\alpha^{k}{(\xi_{+}-\xi^{k})}$ (6a)
@@
where $\alpha^{k}\in(0,1]$ is the chosen step length at iteration $k$. Approximation methods for the Hessian are discussed in Subsection 4.4.1.
3.2 Local Convergence Analysis
@@
The convergence properties of set-valued Newton methods are a long studied subject of variational analysis Dokov1998 ; Dontchev2013 ; mordukhovich_generalized_2021 ; Louzeiro2023 . Local convergence can usually be established provided that the inverse of $(\Phi(\xi^{k},H_{k})+\mathcal{N})$ in (7) is well defined and the approximation $\Phi(\xi^{k},H_{k})$ is ‘sufficiently good’ – the better the approximation, the faster the rate of convergence. In particular, we make the following assumptions.
Assumption 1(Regularity)
The exist constants $\varkappa,\ell_{k}>0$, $k\in\mathbb{N}$, such that
@@
    the inverse $(\Phi(\xi^{k},H_{k})+\mathcal{N})^{-1}$ is locally single-valued around $(\xi_{\star},\lambda_{\star})$ and locally Lipschitz continuous at 0 with constant $\varkappa$ uniformly for $\xi^{k}$ around $\xi_{\star}$ and all $H_{k}$;
@@
    the mapping $\varphi(\xi^{k},\lambda^{k})+\Phi(\xi^{k},H_{k})(\xi^{\prime}-\xi^{k},\lambda^{\prime}-\lambda^{k})$ is locally Lipschitz continuous relative to $(\xi^{k},\lambda^{k})$ at $(\xi_{\star},\lambda_{\star})$ with constant $\ell_{k}$ uniformly for $(\xi^{\prime},\lambda^{\prime})$ around $(\xi_{\star},\lambda_{\star})$ and all $H_{k}$
for all $k\in\mathbb{N}$.
Part (a) of Assumption 1 is called strong regularity of the set-valued mapping $(\Phi(\xi^{k},H_{k})+\mathcal{N})$ and corresponds to unique primal-dual solutions to the convex problem (4) that are continuous under first-order perturbations Dinh2010 . Regularity can be enforced by choosing a positive definite approximation $H_{k}$. The Lipschitz constant in part (b) measures how good the linear operator $\Phi(\xi^{k},H_{k})$ describes the change of $\varphi(\xi^{k},\lambda^{k})$ as $(\xi^{k},\lambda^{k})\to(\xi_{\star},\lambda_{\star})$ and can be thought of as generalization of an error in the derivative of $\varphi$. In particular, Assumption 1(b) is satisfied with $\ell_{k}=0$ if $H_{k}$ is the second-order derivative $\nabla_{\xi\xi}^{2}[f(\xi_{k})-\langle\lambda_{k},g(\xi_{k})\rangle]$.
Theorem 3.1(Local convergence)
@@
Proof
By virtue of (Dontchev2021, , Theorem 8.5) with $h:(\xi,\lambda)\mapsto\varphi(\xi_{\star},\lambda_{\star})+\Phi(\xi_{\star},H_{k})(\xi-\xi_{\star},\lambda-\lambda_{\star})$, the sequence $\{(\xi^{k},\lambda^{k})\}_{k\in\mathbb{N}}$ generated by (7) is unique and satisfies
$\|(\xi^{k+1},\lambda^{k+1})-(\xi_{\star},\lambda_{\star})\|\leq\varkappa\ell_{k}\|(\xi^{k},\lambda^{k})-(\xi_{\star},\lambda_{\star})\|$
for all $k\in\mathbb{N}$, provided that $(\xi^{k},\lambda^{k})$ is close to $(\xi_{\star},\lambda_{\star})$; this implies (super) linear convergence if $\varkappa\ell_{k}<1$ (resp., $\varkappa\ell_{k}\searrow 0$), and quadratic convergence if $\varkappa\ell_{k}\equiv 0$. ∎
3.3 Line Search with Filter
This subsection provides the core components of the filter line search algorithm, which is primarily based on (bieglerNonlinearProgrammingConcepts2010, , p. 120). We outline the core parts to make this paper self-contained. The filter line search algorithm is summarized in Algorithm 2. Three major steps are conducted: 1) check filter acceptance; 2) check sufficient decrease in cost; and 3) augment filter if necessary.
3.3.1 Filter Acceptance
Unlike merit functions, that combine cost $f$ and constraint violation $\theta$ in a single objective, a filter tries to deal with these two conflicting goals in a bi-objective optimization manner. No penalty parameter must be estimated, as is the case for penalty functions or augmented Lagrangian. The filter $\mathcal{F}$ consists of $l\in\mathbb{N}$ so-called dominant pairs $(f(\xi^{l}),\theta(\xi^{l}))\in\mathcal{F}$ which form a forbidden region. In a first step, the current trial point is checked for filter acceptance, that is, checking if the trial point lies not in the forbidden region and makes sufficient progress regarding constraint violation. A new trial point, usually the result of (6) for the candidate step length $\alpha$, is accepted to the filter $\mathcal{F}$ if
$f(\xi^{k+1})<f(\xi^{l})\quad\text{or}\quad\theta(\xi^{k+1})<\theta(\xi^{l}),\quad\text{for all $(f(\xi^{l}),\theta(\xi^{l}))\in\mathcal{F}$}$ (8)
Condition (8) is only an acceptance or rejection criterion, but does not necessarily ensure progress. Therefore, additional sufficient decrease conditions are checked for an acceptable point, as outlined next. If the candidate fails or violates the first condition (8), a second-order correction may be applied, or the step length is reduced.
3.3.2 Sufficient Decrease Conditions
@@
where $s_{\phi},s_{\theta}\in(0,1)$ and $\delta\in(0,1)$ are small constants. This condition ensures a sufficiently large progress for the cost function compared to the current constraint violation bieglerNonlinearProgrammingConcepts2010 . Equations (9) and (10) are also referred as f-type switching condition.
In case the f-type switching is fulfilled and additionally if $\theta(\xi^{k+1})<\theta_{\text{min}}$ wachterImplementationInteriorpointFilter2006a then the Amijo condition is checked
@@
If no reduction in cost is possible (according to f-type) it is checked if either the constraint violation or cost is smaller than the previous iterate with a small envelope, that is,
$\theta(\xi^{k+1})\leq\theta(\xi^{k})(1-\gamma_{\theta})\quad\text{or}\quad\phi(\xi^{k+1})\leq\phi(\xi^{k})-\gamma_{f}\theta(\xi^{k})$ (12)
@@
3.3.3 Filter Augmentation
Unlike fletcherNonlinearProgrammingPenalty2002 , the filter is only augmented if, and only if, at least one of the sufficient decrease conditions (9)–(11) does not hold. This stricter augmentation rule avoids cycling, where the iterates alternately improve one of the two measures but deteriorate the other wachterLineSearchFilter2005 .
3.3.4 Full Algorithm
Algorithm 2 outlines the filter line search procedure (cf.  bieglerNonlinearProgrammingConcepts2010 ). To avoid the Maratos effect nocedalNumericalOptimization2006 , we additionally add a second-order correction (details in Section 4.4.2). If the second-order correction is able to find an acceptable trial point, the line search is terminated, otherwise, the step length is reduced. In case the step-length falls below the minimum step-length $\alpha_{\text{min}}$ a feasibility restoration phase is invoked. The primary goal of the feasibility restoration phase is to minimize the constraint violation, such that the trial point is acceptable to the filter.
@@ 32   Augment filter
Algorithm 2 Filter line search with second-order correction
3.4 Required SOS Modification
The filter frequently needs to check progress in cost and feasibility. While checking progress in cost is a simple function evaluation, evaluating the constraint violation (feasibility) for SOS constraints is more complicated. In parameter optimization, constraint violation can be assessed by evaluating the constraints at the current iterate and use some norm, e.g., $\theta(\xi^{k}):=\|g(\xi^{k})\|_{\infty}$. For conic optimization, the constraint violation check involves a projection onto the corresponding cone (cf. Boyd2005 ). Although some cones possess relatively simple (analytical) projection operators, this is not the case for the SOS cone. Since the constraint violation is frequently evaluated, we are interested in precise yet efficient methods for this task. Similarly, a dedicated feasibility restoration, which primary goal is to reduce constraint violation, must be designed and implemented for SOS.
4 Practical Implementation
This section gives a detailed overview of the implementation of the proposed sequential quadratic SOS algorithm. Algorithm 3 shows the complete optimization procedure. An open-source implementation is available in the recently introduced software CaΣoS Cunis2025a , which is specifically designed to deal with nonconvex SOS problems. Hence, we will shortly outline the core ideas behind CaΣoS. Afterwards, the SOS modifications are discussed.
4.1 Background CaΣoS Framework
The MATLAB-based software CaΣoS was presented in Cunis2025a . Its framework for nonlinear symbolic polynomial expressions is inspired by and built upon CasADi andersson2019 . It implements efficient polynomial data types, linear and nonlinear polynomial operations, and, with the CasADi backend, parameterized solvers and functions. The last feature – novel to sum-of-squares toolboxes – is advantageous for iterative procedures.
@@ 35     Compute Hessian approximation
4.2 Constraint Violation Estimation
@@
We assume $n_{g}$ nonlinear constraints. The constraint violation $\theta(\xi^{k})$ for the solution $\xi^{k}$ indicates whether the constraints $g(\xi^{k})\in\Sigma[x]^{n_{g}}$ are satisfied ($\theta(\xi^{k})=0$) or, if not, measures how far ($\theta(\xi^{k})>0$) the solution is from satisfying the constraints. Take $p_{k}:=g(\xi^{k})$.
4.2.1 SOS Projection
@@
4.2.2 Signed Distance
In (Cunis2025, , Section 4.3), a signed-distance metric for generalized cones is introduced. It utilizes a scalarization function khan2015 to map $p_{k}$ to a finite signed distance between $p$ and the boundary of $\Sigma[x]^{n_{g}}$. The signed-distance metric is defined as
@@ 4.2.2 Signed Distance
$\theta(\xi^{k}):=\begin{cases}0&\text{if }\mathrm{max}(r^{*})\leq\epsilon\\ \mathrm{max}(r^{*})&\text{otherwise}.\end{cases}$ (16)
@@
4.2.3 Pseudo-Projection via Sampling
Verifying that $p_{k}(x)\in\Sigma[x]^{n_{g}}$ can be replaced by checking the necessary condition $p_{k}(x)\geq 0$ on a finite set of sample points $x^{(i)}\in\mathbb{R}^{n}$, $i=1,\dots,M$, and nonnegativity is checked element-wise. Since sampling over $\mathbb{R}^{n}$ is computationally infeasible, the samples are drawn from a bounded hypercube.
@@ 4.2.4 Discussion
Figure 1 compares the computation times of these methods for increasing numbers $n$ of indeterminate variables, using 10 randomly generated SOS polynomials of degree six. The implementation is provided in the supplementary material DARUS-5677_2026 . The projection method scales poorly, becoming impractical for larger $n$. In contrast, the signed distance remains efficient, while sampling with 1,000 points is fastest but least accurate. Based on this comparison, we implement the signed distance method for its strong balance of accuracy and efficiency.
@@
4.3 Feasibility Restoration
Small step lengths can cause the solver to stall or fail. If $\alpha^{k}<\alpha_{\text{min}}$ or infeasibility is detected in (4), Alg. 3 enters the feasibility restoration phase (lines 10 and 21 in Alg. 3) to maintain robustness (nocedalNumericalOptimization2006, , p.439). This phase aims to reduce the constraint violation and provide a solution that is acceptable to the filter of the original problem.
Remark 1
There is no guarantee that the restoration finds asolution acceptable to the filter. If it does not, we want it to converge to a local (infeasible) point. This information might be used by the user to take corrective action on the problem formulation (cf. wachterImplementationInteriorpointFilter2006a ).
The feasibility restoration might be invoked after certain progress has already been made towards a local solution, but the constraint violation is still too high. In that case, we want the restoration not to deviate too much from the current, potentially close to (local) optimal solution. To achieve this, we introduce a weighted regularization term based on the current constraint violation wachterImplementationInteriorpointFilter2006a ; ulbrich2004b .